# Physically Compatible 3D Object Modeling from a Single Image

### v2

@@ 1 Introduction
Related work.
Our method is mainly related to: 1) single-view 3D reconstruction, where our work emphasizes the integration of physical modeling into the reconstruction process; 2) physically based 3D modeling, where we incorporate physical principles as hard constraints within the reconstruction process; and 3) fabrication-aware shape design, where our work directly constructs physical objects from a single input image rather than relying on manual creation of the initial geometry. A comprehensive discussion of related work is provided in Appendix G.
@@ 3.1 Baselines and Evaluation Protocol
Ours 0.45 $\pm$ 0.96 72.4 0.069 $\pm$ 0.048
Ours 1.01 $\pm$ 1.34 85.9 0.116 $\pm$ 0.065
MeshLRM Baseline surface mesh 1.55$\pm$ 2.13 0.54 $\pm$ 1.41 28.6 0.065 $\pm$ 0.042
Ours 0.38 $\pm$ 1.05 73.8 0.064 $\pm$ 0.042
TripoSR Baseline NeRF 1.43 $\pm$ 1.12 0.29 $\pm$ 1.28 24.2 0.066 $\pm$ 0.047
Ours 0.22 $\pm$ 0.94 80.6 0.059 $\pm$ 0.039
Ours 0.19 $\pm$ 0.78 92.4 0.057 $\pm$ 0.040
@@ Appendix F Dynamic Simulation of Deformable Objects
Appendix G Related Work
Single-view 3D reconstruction.
Physics-based 3D modeling.
There has been an increasing interest in incorporating physics into 3D shape modeling. While many approaches utilize video input, which offers a richer temporal context for inferring physical properties such as material parameters [49] and geometry [19], others approach the problem by first reconstructing an object’s geometry from multi-view images and subsequently applying physical simulations [10, 44]. Additionally, several studies have explored extracting physical information from static images [47, 3, 35], using data-driven techniques to estimate properties like shading, mass, and material. In contrast, our work incorporates physical principles, specifically static equilibrium, as hard constraints within the reconstruction process. This integration allows for the optimization of 3D models that adhere to desired physical behaviors depicted by the image.
Fabrication-aware shape design.
Appendix H Limitations and Future Work
Appendix I Broader Impacts

### v3

@@ 1 Introduction
2 Related Work
Single-view 3D reconstruction.
Physics-based 3D modeling.
There has been an increasing interest in incorporating physics into 3D shape modeling. While many approaches utilize video input, which offers a richer temporal context for inferring physical properties such as material parameters [51] and geometry [19], others approach the problem by first reconstructing an object’s geometry from multi-view images and subsequently applying physical simulations [10, 46, 26, 25]. Additionally, several studies have explored extracting physical information from static images [49, 3, 37], using data-driven techniques to estimate properties like shading, mass, and material. In contrast, our work incorporates physical principles, specifically static equilibrium, as hard constraints within the reconstruction process. This integration allows for the optimization of 3D models that adhere to desired physical behaviors depicted by the image.
Fabrication-aware shape design.
@@ 4.1 Baselines and Evaluation Protocol
Ours 0.45 $\pm$ 0.96 72.4 0.069 $\pm$ 0.048
Ours 1.01 $\pm$ 1.34 85.5 0.116 $\pm$ 0.065
MeshLRM Baseline surface mesh 1.55$\pm$ 2.13 0.54 $\pm$ 1.41 29.6 0.065 $\pm$ 0.042
Ours 0.38 $\pm$ 1.05 74.5 0.064 $\pm$ 0.042
TripoSR Baseline NeRF 1.43 $\pm$ 1.12 0.29 $\pm$ 1.28 24.2 0.066 $\pm$ 0.047
Ours 0.22 $\pm$ 0.94 80.6 0.059 $\pm$ 0.039
Ours 0.19 $\pm$ 0.78 92.2 0.057 $\pm$ 0.040
@@ Appendix F Dynamic Simulation of Deformable Objects
Appendix G Broader Impacts